Based on the following reaction; Cl2(g) + NaClO(aq) → NaCl(aq) + ClO2(g) (Not balanced) 2.5 L of Cl2(g) measured at 20.0 °C and 4.12 atm is reacted with 800 mL of 2.30 M NaClO(aq). If 51.8 g of ClO2(g) is obtained, what is the percent yield of the reaction? Determine the limiting and remaining (as gram) reagent? 


1
Expert's answer
2021-11-24T12:21:59-0500

Molar mass of NaClO= 74.44

2.3×74.44=171.212

=800/171.212=4.67

= 4.67×74.44=347.83

347.83/800×100

=43.48%